OpenStudy (anonymous):

How many solutions does the absolute value |x-5|=4 have?

OpenStudy (solomonzelman):

When you are solving this type of equation the first step is, |x-5|=4 split into 2 different equation x-5=4 or x-5= -4

OpenStudy (solomonzelman):

So you will have 2 solutions.

OpenStudy (solomonzelman):

If you have any questions, please do NOT hesitate to ask.